The ISS web page for Perdigao is now up and includes links to quick look summary plots from the 1290 MHz wind profiler site. It is available at :
Thanks to Gary for setting up the data transfer and summary plots.
The WXT surface met sensor stopped reporting for an as yet undetermined reason. Attempted to replace with another WXT on the same DSM port but that gave the same sensor not found error, suggesting a DSM problem. To replace it we installed a Lufft ws700 adjacent to the WXT and that appears to be recognised by the DSM and it's data is being ingested (from around 11UTC, 17 Dec).
Weather mainly cloudy with cold northerly breeze, rain by evening. All systems are working well. We tidied up around the sites, repaired a hole in the roof of the profiler site shelter and installed a vent fan to prevent overheating. Replaced webcam since the previous one had failed due to a rain. Put the DSM at the sodar site and the cell modem at the profiler site on UPS. Continued surveying possible sites for the NCAS wind profiler.
This was the last day of ISS setup. Jen and Gary have already left; Lou, John and Bill are leaving tomorrow - many thanks to all!
ISS is running with setup largely complete - the last major item was installation of acoustic shielding surrounds for the wind profiler RASS, which were installed today. RASS was turned on this afternoon, as usual running for 5 minutes every half hour. It is currently measuring virtual temperature up to about 600 - 700m altitude, with the amp running at moderate power - about 7 VAC (amp output), which produces around 115 dB at the speakers (50 dB at 30m horizontally out from the shielded speakers). Winds are being measured up to the 2 - 3km level.
The cell modem at the sodar site was switched to a SIM card from INEGI (still on the NOS provider, but under a long term contract rather than the 15 day service of the previous card). The profiler site continues to use a Vodafone mobile HotSpot which should be active until about Jan 20 by which time the valley wireless network should be up and ready to take over.
We visited potential sites for the NCAS wind profiler east of the valley, and identified a suitable area
Weather was cool and cloudy with light winds, switching to showers by evening.
Today was rainy, with heavy falls early, gradually easing through the day to occasional scattered showers by mid afternoon. It was also Jen, Gary and Eckhard's last day, many thanks for their efforts. Gary got network access working at both sites (via cell modems); although they are not yet transferring data automatically, it is possible to remotely access the two sites. Eckhard did more checks on the sodar-RASS, it is working very well and seems to be unaffected by the nearby high tension power lines. The wind profiler is also working well with reduced clutter since installing the clutter screen. The heavy rain revealed a couple of minor problems, a small leak at the profiler shelter and an intermittent connection to the webcam, both of which will be repaired as conditions dry out.
Work continued at both the sodar and the profiler sites. At the profiler site, we installed the clutter fence on the wind profiler. Although I've only had a chance to examine a short period of data afterwards, there did appear to be a significant improvement with the wind turbine being less obvious in the signal plots (the fence was installed 14:30 - 15:30 UTC). At the sodar site, power was finally installed this afternoon and Eckhard got the sodar working soon afterwards. It is working very well, reporting winds to the 300 - 400 meter range, and RASS temperature to almost as high. Gary worked on network connections at both sites, although had limited success with various cell modems and router combinations. It appears we need to purchase a WiFi hotspot for the profiler site, but may get our NOS USB modem working at the sodar site. Today the weather was cold, with heavy dense fog all morning, particularly in the valley.
Today we split into two groups: Lou, Gary and I worked at the profiler site, while Jen and John helped Eckhard at the sodar site. The sodar is now assembled although power won't be available to test the system at that site until tomorrow. At the profiler site, further configuration work was done such as various updates on the DM computer, web plots, adjustments to the WXT and webcam, as well as more sorting of equipment at the seatainer. This morning there was dense fog over the profiler site until midday, although the sodar site was mainly clear. Winds were very light.
ISS setup continues. Transported more sodar equipment to to the farm site, cleared weeds at that site, and organized crates and the seatrainer. Eckhard Gast arrived from Metek in Germany and will begin setting up the sodar-RASS tomorrow. Gary worked on the DM computer, so now it is ingesting profiler and WXT data, plus webcam images. The profiler running well - seeing a little clutter from the wind turbine although NIMA is did a good job of filtering that in todays light winds. Winds consistently to 2km and often to 3km. Sunny with very light winds.
We ran power lines from the power drop at the MEO cell tower, completed the shelter, assembled and connected the wind profiler electronics, computer and RASS. Turned on the profiler for a test run for about an hour, including a brief RASS run - it appeared to be working well with winds up to about 1 km and RASS to about 600m. However there was lots of ground clutter because we haven't installed the clutter screen yet, nor the RASS surrounds. Plan more tests tomorrow. We also installed the WXT and webcam on the tower, although haven't connected those yet. Weather cloudy this morning, then clearing with light variable winds.
Today we started work on the profiler antenna, erected a 3m tower, and continued moving equipment and worked on the shelter. In addition power hookup was installed at the site (coming from the adjacent cell tower). The weather continues to cooperate, being clear with light winds and a high around 15C.
ISS Setup is getting underway after a short delay. The past few days have been organizing and receiving equipment, meeting with locals including a group from IPMA in Lisbon and INERGI in Porto. The profiler site was leveled on Wednesday by the local municipality, although power is not yet available there or at the sodar site. Our seatainer sits at an olive processing shed at the foot of the access road to the profiler site and we have been gradually transporting equipment up to the site. (The access road is too rough and steep for our usual procedure of locating the seatainer at the site and using as office space). The municipality helped us move some larger items (such as the profiler) with one of their trucks Tuesday, but we are moving most equipment with our small rental pickup.
The sodar arrived from Metek in Germany on Tuesday a couple of days ahead of schedule and was delivered to the Ops Center, although was in a larger and heavier set of crates than expected. We started unpacking and transporting equipment to the sodar site and started clearing that site of bushes and weeds.
Today the main activity was building an equipment shelter at the profiler site.